water rises in a capilary tube of certain height such that the upward force due to surface tension is balanced by 75*10^-4N force due to the weight of the liquid If the surface tension of water is 6*10^-2 N/m the inner circumference of the capillary must be

Grade:12

1 Answers

Sudheesh Singanamalla
13 years ago

weight of liquid = upward force due to surface tension

75 * 10^-4 = 2*pi*r*T

circumference = 2*pi*r = (75*10^-4)/(6*10^-2) =0.125 = 12.5* 10^-2m

 

Please approve if this is the right answer